# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 9.2.1 Brero, Eden, Gerstgrasser, Parkes, and Rheingans-Yoo (2021): RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).146146146A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).147147147In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),148148148Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.
@@
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,149149149The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.150150150DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 9.3 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)151151151DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.
@@ 10.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.152152152Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
@@ 10.1.2 Minimax-Q Learning
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.153153153The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.154154154In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
@@ 10.1.3 Nash-Q Learning
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.155155155In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
@@ 10.1.4 The Convergence Problem
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.156156156Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
@@ 10.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.157157157Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
@@ 10.2 Counterfactual Regret Minimization
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:158158158Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).159159159An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).160160160The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).

### 17 Mar 2026 v3

@@ 8.2.1 Brero, Eden, Gerstgrasser, Parkes, and Rheingans-Yoo (2021): RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).153153153A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).154154154In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),155155155Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.
@@
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,156156156The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.157157157DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 8.3 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)158158158DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. Moll (2025) argues that rational expectations equilibria in heterogeneous agent models are computationally intractable and proposes reinforcement learning as a more tractable alternative for modeling how agents form beliefs and make decisions; see also the textbook treatment in Zhao (2025). This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.
@@ 8.4 Optimal Policy Design
8.5 Simulation Study: DDC Estimation at Scale
The test bed is a multi-component extension of the Rust (1987) bus engine replacement model. In the original formulation a maintenance superintendent observes discretized mileage $m\in\{0,\ldots,M{-}1\}$ and makes a binary keep-or-replace decision, facing running cost $c(s;\theta)=\theta_{1}x+\theta_{2}x^{2}$ with $x=m/M$, replacement cost $RC$, and Type I extreme value additive errors that yield logit conditional choice probabilities. We extend the model to $K$ independent wear components, each evolving in $\{0,\ldots,M{-}1\}$, with aggregate normalized wear $x(s)=\sum_{k}m_{k}/M$ entering the same cost function. The state space is $|\mathcal{S}|=M^{K}$, so increasing $K$ produces a controlled scaling experiment in which the data-generating process is identical across scales but the computational burden grows exponentially.
We compare four estimation methods on a multi-component bus engine replacement problem (Rust, 1987) with $K\in\{1,2,3,4\}$ independent wear components, producing state spaces from 20 to 160,000 states. Panel data consist of $N=500$ agents observed for $T=100$ periods. The four methods are NFXP (nested fixed-point MLE), CCP (Hotz-Miller inversion), TD-CCP Linear (semi-gradient TD with polynomial basis), and TD-CCP Neural (approximate value iteration with a two-layer MLP), where the two TD-CCP variants follow Adusumilli et al. (2022). Each configuration is replicated across 5 seeds; Table 15 and Figure 15 report means.
Table 15: DDC estimation results across state-space scales. Rows show method, number of components $K$, state-space size $|\mathcal{S}|$, mean wall-clock time (seconds), RC bias, and root mean squared error for each structural parameter. Averages over 5 seeds; dashes indicate method failure (sparse state coverage).
Method K $|S|$ Time (s) RC Bias RC RMSE $\theta_{1}$ RMSE $\theta_{2}$ RMSE
NFXP 1 20 0.2 +0.074 0.102 0.363 0.618
CCP 1 20 0.1 +0.076 0.103 0.374 0.639
TD-CCP Linear 1 20 0.1 -0.083 0.096 0.178 0.222
TD-CCP Neural 1 20 33.7 +0.118 0.133 0.487 0.819
NFXP 2 400 0.4 -0.023 0.077 0.218 0.290
CCP 2 400 0.1 +0.042 0.085 0.461 0.775
TD-CCP Linear 2 400 0.1 -0.211 0.217 0.131 0.844
TD-CCP Neural 2 400 39.3 -0.040 0.070 0.099 0.168
NFXP 3 8,000 4.2 +0.004 0.043 0.290 0.454
CCP 3 8,000
TD-CCP Linear 3 8,000 0.1 -0.251 0.253 0.163 1.171
TD-CCP Neural 3 8,000 39.2 -0.009 0.046 0.266 0.417
NFXP 4 160,000 172.6 -0.036 0.070 0.130 0.113
CCP 4 160,000
TD-CCP Linear 4 160,000 0.2 -0.333 0.337 0.217 1.119
TD-CCP Neural 4 160,000 44.1 -0.029 0.069 0.167 0.144
Figure 15: Wall-clock estimation time versus state-space scale for the four DDC estimators. Vertical axis is log-scaled. Each point is the mean over 5 seeds.
NFXP scales from 0.2s at K=1K{=}1 ($|\mathcal{S}|{=}20$) to 179s at K=4K{=}4 ($|\mathcal{S}|{=}160{,}000$), reflecting the cost of repeated value iteration inside each likelihood evaluation. CCP becomes infeasible beyond K=2K{=}2 due to sparse state coverage in the panel data. TD-CCP Neural maintains near-constant runtime (${\sim}28$$44$s) across all $K$ levels with competitive accuracy (RC RMSE 0.077 at K=4K{=}4), validating the AVI approach of Adusumilli et al. (2022). TD-CCP Linear runs in under 0.3s at all scales but suffers from basis misspecification at higher $K$ (RC RMSE 0.337 at K=4K{=}4); see Table 15 and Figure 15.
@@ 9.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.159159159Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
@@ 9.1.2 Minimax-Q Learning
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.160160160The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.161161161In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
@@ 9.1.3 Nash-Q Learning
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.162162162In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
@@ 9.1.4 The Convergence Problem
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.163163163Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
@@ 9.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.164164164Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
@@ 9.2 Counterfactual Regret Minimization
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:165165165Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).166166166An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).167167167The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).